# Black Hole States in Quantum Spin Chains

### v1

@@ Abstract
We define a black hole state in a spin chain by studying thermal correlators in holography. Focusing on the Heisenberg model we investigate the thermal and complexity properties of the black hole state by evaluating its entanglement entropy, emptiness formation probability and Krylov complexity. The entanglement entropy grows logarithmically with effective central charge $c\simeq 5.2$. We find evidence for thermalization at infinite temperature.
@@ I The Black Hole State
At the heart of the gauge–gravity duality lies a remarkable identification of a thermal ensemble with a black hole in the dual gravitational description [1]. The infall of a particle into the black hole gives rise to thermal one-point functions for local operators. From the gravity point of view, the one-point functions serve as sensitive probes of the black-hole interior [2].
In a more refined holographic picture particles are replaced by strings, which in turn admit an effective spin-chain description. The key advantage of the latter is manifest integrability [3]. The one-point functions (in other setups where they arise) map to overlaps with the spin-chain boundary states and can be efficiently studied by integrability methods [4, 5] (see [6, 7] for a review). Since black holes are chaotic systems, we expect the integrability to be broken in the thermal ensemble, and we will later confirm this expectation.
@@
We will construct what we call the black-hole states in this model using simple plausibility arguments. Their direct counterparts govern thermal one-point functions in AdS/CFT as we show later.
@@
The singlet formed by two spins is their anti-symmetric combination: $\varepsilon_{\alpha_{1}\alpha_{2}}$, where $\alpha_{i}\in\{\uparrow,\downarrow\}$. Four spins can be split in two pairs in three possible ways, but only two of them lead to linearly independent singlet states, because of the identity $\varepsilon_{\alpha_{1}\alpha_{2}}\varepsilon_{\alpha_{3}\alpha_{4}}+\varepsilon_{\alpha_{1}\alpha_{4}}\varepsilon_{\alpha_{2}\alpha_{3}}=\varepsilon_{\alpha_{1}\alpha_{3}}\varepsilon_{\alpha_{2}\alpha_{4}}$ [8], illustrated in fig.1a. Rumer’s identity can be used to eliminate crossings for any number of spins by expressing them through non-crossing (”kreuzungslose”) pairings. Those are linearly independent and form a comlete albeit non-orthogonal basis [9]. The number of pairings, for $L$ spins, equals the Catalan number $C_{L/2}$, and this is precisely the dimension of the singlet subspace [10].
We define the black-hole state as the sum of all non-crossing pairings taken with equal weights. The state can be built recursively from
@@
as illustrated in fig. 1. Catalan numbers, let us mention in passing, satisfy pictorially equivalent recursion relation.
The non-crossing pairings are identical to t’ Hooft’s planar diagrams, and the wavefunction of the black-hole state can be alternatively defined by a matrix integral:
@@
The Heisenberg model is integrable as its Hamiltonian commutes with an infinte number of conserved charges, starting with $Q_{3}=i\sum_{\ell}\bm{\sigma}_{\ell}\cdot[\bm{\sigma}_{\ell+1}\times\bm{\sigma}_{\ell+2}]$. A boundary state preserves this structure if annihilated by $Q_{3}$ (and higher odd charges as well) [12, 13]. Integrable boundary states have remarkable mathematical properties, in particular, their overlaps with the physical eigenstates of the Hamiltonian can be systematically constructed by Bethe ansatz [14, 15, 16]. The black-hole state however breaks integrability. It is easy to check that $Q_{3}\left|{\rm BH}\right\rangle\neq 0$ starting at length six and higher.
We will probe the black-hole state with the standard diagnostics of quantum chaos: entanglement entropy, Krylov complexity and eigenvalue thermalization. Despite integrability of the underlying Hamiltonian, we will find features of typical chaotic behavior, such as the logarithmic growth of entanglement entropy or thermalization at infinite temperature.
To contrast this behavior with the regularity characteristic of integrable models, we consider a singlet Valence-bond state (VBS):
${\rm VBS}^{\alpha_{1}\ldots\alpha_{L}}=\varepsilon^{\alpha_{1}\alpha_{2}}\ldots\varepsilon^{\alpha_{L-1}\alpha_{L}},$ (5)
which is known to preserve integrability [17]. It has the same quantum numbers as the black-hole state upon averaging over translations by one site: $\left\langle{\rm VBS}_{\pi}\right|=\left\langle{\rm VBS}\right|(1-T),$ but has very different entanglement properties and much lower Krylov complexity.
@@ II Entanglement Entropy
where $c$ is the central charge of the CFT and $c_{1}^{\prime}$ is a non-universal constant. This is in contrast to a gapped ground state which will have constant entanglement entropy and thermal states for which the entanglement entropy grows linearly with $\ell$ [20]. The behaviour in (12) was confirmed for the ground state of the critical Heisenberg XXX1/2 chain [21, 22, 23]. The central charge in this case equals one: $c=1$.
@@ III Thermalization
One way to formulate the Eigenvalue Thermalization Hypotesis (ETH) posits that time averages in a sufficiently generic state thermalize [24, 25]:
@@
Course-graining in the infinite-volume limit converts, it is believed, the diagonal ensemble into the grand-canonical one provided the operator acts locally and the initial state is sufficiently generic [26]. The grand-canonical ensemble of an integrable model contains infinitely many chemical potentials, one for each conserved charge, and ETH does not imply such a remarkable reduction in the degrees of freedom as for chaotic systems. Surprisingly, the black-hole state seems to thermalize without any dependence on the higher conserved charges, not even on the Hamiltonian itself.
We conjecture that the course-grained density matrix of the diagonal ensemble is the projector onto the singlet subspace:
@@
where $dg$ is the Haar measure on the $SU(2)$ group manifold. The black-hole state thermalizes into the Gibbs ensemble at infinite temperature.
To test the conjecture we study the emptiness formation probability (EFP) [27]:
@@
Its expectation value describes a probability to find $n$ consecutive spins in the up state and decays exponentially with the length of the string: $\left\langle P(n)\right\rangle\sim\,{\rm e}\,^{-fn}$ with the exponent given by the free energy density [28, 29, 30]. The latter is temperature-dependent, making EFP an ideal probe of thermalization.
@@
It is easy to calculate the EFP in the infinite-temperature ensemble (16). In the parametrization $g=n_{0}+i\sigma^{i}n_{i}$ with a unit four-vector $n_{\mu}$ the group integrals become averages over the round sphere, for instance:
@@
reconfirming that the total number of singlets is the Catalan number. Here we have taken into account that $\mathop{\mathrm{tr}}g=2n_{0}$. By the same token, $\mathop{\mathrm{tr}}g(1+\sigma^{3})/2=n_{0}+in_{3}$, and consequently
@@
In fig. 4 we compare this with the numerical results for the diagonal ensemble associated with the black-hole state, for spin chains of length from $8$ to $16$. The results agree reasonably well. A certain scatter in the plot can be attributed to finite-size corrections which are appreciable for relatively short spin chains accessible to direct diagonalization.
@@ IV Krylov Complexity
Krylov complexity measures to what extent an operator or a state explores a system’s Hilbert space under time evolution, i.e. when repeatedly acted upon by the Hamiltonian  [31, 32]. An extensive recent review of the virtues and applicability of the concept can be found in [33]. The Krylov complexity of a state, also known as the spread complexity, depends both on the Hamiltonian and on the state itself [34, 35, 36, 33, 37]. We will be interested in studying the Krylov complexity of the black hole state (2). For comparison we also calculated the Krylov complexity of the integrable valence bond state (5).
@@
A more elaborate version can be found in [32]. To get rid of short-scale oscillations one often invokes time averaging:
@@
In fig. 5 we show the time averaged Krylov complexity of the black hole state as well as the integrable two-site singlet, for $L=16$.
@@ V Thermal One-point Functions
We finally return to our original motivation of studying thermal one-point functions. To this end, we consider scalar operators in the $\mathcal{N}=4$ super-Yang-Mills theory:
@@
where $i_{1},\ldots,i_{L}\in\{1,\ldots,6\}$. The one-loop dilatation operator operator in this sector is the Hamiltonian of an integrable $SO(6)$ spin chain: $H=\sum_{\ell}h_{\ell,\ell+1}$ with h=22P+Kh=2-2P+K, where $P$ and $K$ are the permutation and trace operators [3].
To compute the one-point function at leading order, we just need to contract identical fields pairwise with the thermal propagator
$D_{\beta}(x)=\frac{1}{4\pi^{2}}\sum_{n=-\infty}^{+\infty}\frac{1}{\left(t+n\beta\right)^{2}+\mathbf{x}^{2}}\,.$ (30)
Thermal effects are encoded in the difference $D_{\beta}-D_{\infty}$, which amounts in omitting the $n=0$ term. Each bubble then contributes
@@
The planar contribution is accounted for by the following spin chain overlap:
$\left\langle\mathcal{O}\right\rangle=\frac{1}{\sqrt{L}}\left(\frac{\pi T}{\sqrt{3\lambda}}\right)^{L}\frac{\left\langle{\rm BH}\right.\!\left|\Psi\right\rangle}{\left\langle\Psi\right.\!\left|\Psi\right\rangle^{\frac{1}{2}}}\,$, (32)
where the boundary state is a direct counterpart of the black-hole state that we have studied in the Heisenberg model:
${\rm BH}^{i_{1}\ldots i_{L}}=\lim_{N\rightarrow\infty}\int d\phi^{i}\,\,{\rm e}\,^{-\,\mathop{\mathrm{tr}}\phi^{i}\phi^{i}}\mathop{\mathrm{tr}}\phi^{i_{1}}\ldots\phi^{i_{L}},$ (33)
with the integration being over six Hermitian $N\times N$ matrices. The state can be also obtained from a recursion relation:
@@
This state shares many features its $SU(2)$ counterpart. In particular, it breaks integrability, as $Q_{3}\left|{\rm BH}\right\rangle\neq 0$ for $Q_{3}=i\sum_{\ell}[h_{\ell,\ell+1},h_{\ell+1,\ell+2}]$.
@@
We have introduced the concept of a black hole state in a quantum spin chain motivated by thermal correlators in holography. As an initial step we concentrated on the black-hole states in the Heisenberg spin chain and demonstrated that they have some characteristics of chaotic nature. One interesting feature is logarithmic growth of entanglement entropy typical for a critical, gapless system.
A black hole state which arises directly from thermal correlators in holography is also not integrable. We leave a comprehensive investigation of its thermal and complexity properties to future work. Besides giving information about the infall of a particle into the black hole [2] and thermal phase transitions [1], thermal one-point functions have an important role to play as input to the thermal bootstrap program [38, 39, 40], not least in relation to the investigation of two-point correlators as probes sensitive to the internal structure of the space-time singularity [41, 42, 43].

### v2

@@ Abstract
We define a black hole state in a spin chain by studying thermal correlators in holography. Focusing on the Heisenberg model we investigate the thermal and complexity properties of the black hole state by evaluating its entanglement entropy, emptiness formation probability and Krylov complexity. The entanglement entropy grows logarithmically with effective central charge $c\simeq 5.2$. We also find evidence for thermalization at infinite temperature.
@@ I The Black Hole State
At the heart of the gauge–gravity duality lies a remarkable identification of a thermal ensemble in field theory with a black hole in the dual gravitational description [1]. The infall of a particle into the black hole permits local operators to exist in isolation and gives rise to their thermal one-point functions which thereby serve as sensitive probes of the black-hole interior [2].
In a more refined holographic picture particles are replaced by strings, which in turn admit an effective spin-chain description. The key advantage of the latter is manifest integrability [3]. The one-point functions (in other setups where they arise) map to overlaps of the spin-chain boundary states and can be efficiently studied by integrability methods [4, 5] (see [6, 7] for a review). Since black holes are chaotic systems, we expect the integrability to be broken in the thermal ensemble, and we will later confirm this expectation.
@@
We will construct what we call the black-hole states in this model using simple plausibility arguments. We show later that their direct counterparts govern thermal one-point functions in AdS/CFT.
Other approaches to defining black hole states based on discrete models and thermal correlation functions can be found in [8], where the black hole emerges from the 2D Ising model, and in [9, 10, 11] where it occurs through a discretization of hyperbolic space.
@@
The singlet formed by two spins is their anti-symmetric combination: $\varepsilon_{\alpha_{1}\alpha_{2}}$, where $\alpha_{i}\in\{\uparrow,\downarrow\}$. Four spins can be split in pairs in three possible ways, but only two of them lead to linearly independent singlet states when anti-symmetrized, because of the identity $\varepsilon_{\alpha_{1}\alpha_{2}}\varepsilon_{\alpha_{3}\alpha_{4}}+\varepsilon_{\alpha_{1}\alpha_{4}}\varepsilon_{\alpha_{2}\alpha_{3}}=\varepsilon_{\alpha_{1}\alpha_{3}}\varepsilon_{\alpha_{2}\alpha_{4}}$ [12] illustrated in fig.1a. Rumer’s identity can be used to eliminate crossings for any number of spins. The non-crossing (”kreuzungslose”) pairings are linearly independent and form a complete albeit non-orthogonal basis of singlets [13]. The number of pairings, for $L$ spins, is the Catalan number $C_{L/2}$, and this is precisely the dimension of the singlet subspace in the spin chain of length $L$ [14].
We define the black-hole state as the sum of all non-crossing pairings taken with equal weights. The state can be built recursively from the relation
@@
as illustrated in fig. 1. Catalan numbers, let us mention in passing, satisfy pictorially equivalent recursion formula.
The non-crossing pairings are identical to t’ Hooft’s planar diagrams, and the wavefunction of the black-hole state can be alternatively computed from a matrix integral:
@@
The Heisenberg model is integrable, its Hamiltonian commutes with an infinite number of extra conserved charges starting with $Q_{3}=i\sum_{\ell}\bm{\sigma}_{\ell}\cdot[\bm{\sigma}_{\ell+1}\times\bm{\sigma}_{\ell+2}]$. A boundary state is compatible with this structure if annihilated by $Q_{3}$ (and higher odd charges as well) [16, 17]. Integrable boundary states have remarkable mathematical properties, in particular, their overlaps with the physical eigenstates of the Hamiltonian can be systematically computed by Bethe ansatz [18, 19, 20]. The black-hole state however breaks integrability. It is easy to check that $Q_{3}\left|{\rm BH}\right\rangle\neq 0$ at length six and higher. We therefore expect that the black hole state exhibits some chaotic features even if the underlying Hamiltonian is completely integrable.
We will probe the black-hole state with the standard diagnostics of quantum chaos: entanglement entropy, Krylov complexity and eigenvalue thermalization. For comparison, we also consider a singlet valence-bond state (VBS) which is known to preserve integrability [21]:
${\rm VBS}^{\alpha_{1}\ldots\alpha_{L}}=\varepsilon^{\alpha_{1}\alpha_{2}}\ldots\varepsilon^{\alpha_{L-1}\alpha_{L}}.$ (5)
It has the same quantum numbers as the black-hole state upon averaging over translations by one site: $\left\langle{\rm VBS}_{\pi}\right|=\left\langle{\rm VBS}\right|(1-T)$. As we shall see it has very different entanglement properties and much lower Krylov complexity.
@@ II Entanglement Entropy
where $c$ is the central charge of the CFT and $c_{1}^{\prime}$ is a non-universal numerical constant. This is in contrast to a gapped ground state which will have constant entanglement entropy and thermal states for which the entanglement entropy grows linearly with $\ell$ [24]. The behavior in (12) was confirmed for the ground state of the critical Heisenberg XXX1/2 chain [25, 26, 27]. The central charge in this case equals one: $c=1$.
@@ III Thermalization
One possible formulation of the Eigenvalue Thermalization Hypothesis (ETH) posits that time averages in a sufficiently generic state thermalize [28, 29, 30]:
@@
Of course the diagonal ensemble does not literally coincide with the grand-canonical one, but course-graining in the infinite-volume limit converts, it is believed, one to the other provided the initial state $\left|\Psi\right\rangle$ is sufficiently generic [31]. The grand-canonical ensemble of an integrable model contains infinitely many chemical potentials, one for each conserved charge, and ETH does not imply such a remarkable reduction in the degrees of freedom as for chaotic systems. Surprisingly, the black-hole state seems to thermalize without any dependence on the higher conserved charges, not even on the Hamiltonian itself.
We find evidence for the course-grained density matrix of the diagonal ensemble being the projector onto the singlet subspace:
@@
where $dg$ is the Haar measure on the $SU(2)$ group manifold. In this sense the black-hole state thermalizes at infinite temperature. The density matrix then is the unit operator but the black-hole state lies in the singlet subspace and therefore the density matrix is a projector.
To probe the diagonal ensemble of the black-hole state we study the emptiness formation probability (EFP) [32]:
@@
that measures a chance of finding $n$ consecutive spins in the up state. The expectation value of EFP decays exponentially with the length of the string: $\left\langle P(n)\right\rangle\sim\,{\rm e}\,^{-fn}$ where the exponent is the free energy density [33, 34, 35]. The latter is temperature-dependent, making EFP an ideal probe of thermalization.
@@
It is easy to calculate the EFP in the infinite-temperature ensemble (16). In the parametrization $g=n_{0}+i\sigma^{i}n_{i}$ with a unit four-vector $n_{\mu}$ the group integrals become averages over the round three-dimensional sphere, for instance:
@@
reconfirming that the total number of singlets is the Catalan number. Here we have taken into account that $\mathop{\mathrm{tr}}g=2n_{0}$. By the same token, $\mathop{\mathrm{tr}}g(1+\sigma^{3})/2=n_{0}+in_{3}$, and the EFP in the infinite-temperature ensemble is
@@
In fig. 4 we compare this prediction with the numerical results for the diagonal ensemble of the black-hole state, for spin chains of length $8$ to $16$. The results agree reasonably well. A certain scatter in the plot can be attributed to finite-size corrections which are appreciable for relatively short spin chains accessible to direct diagonalization.
@@ IV Krylov Complexity
Krylov complexity measures to what extent an operator or a state explores a system’s Hilbert space under time evolution, i.e. when repeatedly acted upon by the Hamiltonian  [36, 37]. An extensive recent review of the virtues and applicability of the concept can be found in [38, 39]. The Krylov complexity of a state, also known as the spread complexity, depends both on the Hamiltonian and on the state itself [40, 41, 42, 43, 39, 44]. We will be interested in studying the Krylov complexity of the black hole state (2). For comparison we also calculated the Krylov complexity of the integrable valence bond state (5). Krylov complexity of coherent states of relevance for the semi-classical limit of AdS/CFT integrable models was discussed in [45]. Furthermore, holographic Krylov complexity was argued to be directly extractable from string motion in AdS [46].
@@
A more elaborate version can be found in [37]. Time averaging is often invoked to get rid of short-scale oscillations:
@@
In fig. 5 we show the time averaged Krylov complexity of the black hole state compared to that of the integrable two-site singlet, for $L=16$.
@@ V Thermal One-point Functions
We finally return to our original motivation of studying thermal one-point functions in AdS/CFT. We consider the full set of scalar operators in the $\mathcal{N}=4$ super-Yang-Mills theory:
@@
where $i_{1},\ldots,i_{L}\in\{1,\ldots,6\}$. The one-loop dilatation operator acting on these states can be identified with the Hamiltonian of an $SO(6)$ spin chain: $H=\sum_{\ell}h_{\ell,\ell+1}$ with h=22P+Kh=2-2P+K, where $P$ and $K$ are the permutation and trace operators [3], and this Hamiltonian is integrable.
To compute the one-point functions in the thermal ensemble, to the leading order in perturbation theory, we need to contract identical fields pairwise with the thermal propagator
$D_{\beta}(x)=\frac{1}{4\pi^{2}}\sum_{n=-\infty}^{+\infty}\frac{1}{\left(t+n\beta\right)^{2}+\mathbf{x}^{2}}\,$, (30)
evaluated at coincident points. Thermal effects are encoded in the difference $D_{\beta}-D_{\infty}$, which amounts in omitting the $n=0$ term. Each bubble then contributes
@@
The full planar contribution to the one-point function is accounted for by the following spin chain overlap:
$\left\langle\mathcal{O}\right\rangle=\frac{1}{\sqrt{L}}\left(\frac{\pi T}{\sqrt{3}}\right)^{L}\frac{\left\langle{\rm BH}\right.\!\left|\Psi\right\rangle}{\left\langle\Psi\right.\!\left|\Psi\right\rangle^{\frac{1}{2}}}\,$, (32)
where the prefactor accounts for the unit normalization of the two-point function. The boundary state is the direct counterpart of the black-hole state that we have studied in the Heisenberg model:
${\rm BH}^{i_{1}\ldots i_{L}}=\lim_{N\rightarrow\infty}\int d\phi^{i}\,\,{\rm e}\,^{-\,\mathop{\mathrm{tr}}\phi^{i}\phi^{i}}\mathop{\mathrm{tr}}\phi^{i_{1}}\ldots\phi^{i_{L}}.$ (33)
The integration here is over six Hermitian $N\times N$ matrices with the usual numerical entries. The state can be also constructed from a recursion relation:
@@
This state shares many features with its $SU(2)$ counterpart. In particular, it breaks integrability: one can check that $Q_{3}\left|{\rm BH}\right\rangle\neq 0$ for $Q_{3}=i\sum_{\ell}[h_{\ell,\ell+1},h_{\ell+1,\ell+2}]$. In contradistinction to the Heisenberg model, the planar contractions do not constitute a full basis of singlets for the $SO(6)$ spin chain, as Rumer identity does not hold for $SO(6)$ spins.
@@
We have introduced the concept of a black-hole state in a quantum spin chain motivated by thermal correlators in holography. As an initial step we concentrated on the black-hole states in the Heisenberg spin chain and demonstrated that they have some characteristics of chaotic nature. One interesting feature is logarithmic growth of entanglement entropy typical for a critical, gapless system, for which we do not have a systematic explanation.
A black-hole state which arises directly from thermal correlators in holography is also not integrable. We leave a comprehensive investigation of its thermal and complexity properties to future work. Besides giving information about the infall of a particle into the black hole [2] and thermal phase transitions [1], thermal one-point functions have an important role to play as input to the thermal bootstrap program [47, 48, 49], not least in relation to the investigation of two-point correlators as probes sensitive to the internal structure of the space-time singularity [50, 51, 52].